OpenStudy (anonymous):

Help! Write the equation of a line that passes through the point (-2, 1) and has a slope of 4.

jimthompson5910 (jim_thompson5910):

y = mx + b y = 4x + b ... plug in the given slope m = 4 1 = 4(-2) + b ... plug in the given point 1 = -8 + b ... ... b = ???

OpenStudy (anonymous):

Use on point slope form: \[y - y_1 = m(x - x_1)\] here m = 4, \(x_1\) = -2 and \(y_1\) = 1